verify the function u(x,y)=e^xsiny is one of the solution of uxx+uyy=0



1
Expert's answer
2022-02-07T17:25:40-0500

"\displaystyle
u=u(x,y)=e^x\sin y\\"

Now,

"\displaystyle
\frac{\partial u}{\partial x}=u_x=e^x\sin y,\ \frac{\partial^2u}{\partial x^2}=u_{xx}=e^x\sin y\\
\frac{\partial u}{\partial y}=u_y=e^x\cos y,\ \frac{\partial^2u}{\partial y^2}=u_{yy}=-e^x\sin y"

"\displaystyle
u_{xx}+u_{yy}=e^x\sin y+(-e^x\sin y)=e^x\sin y-e^x\sin y=0"

Thus, "\displaystyle
u=u(x,y)=e^x\sin y\\" is a solution of the PDE "u_{xx}+u_{yy}=0"